OpenStudy (anonymous):

Suppose you invest $500 at an annual interest rate of 8.2% compounded continuously. How much will you have in the account after 15 years?

OpenStudy (anonymous):

@Schrodingers_Cat can you help me again?

OpenStudy (nurali):

A = 500 e^(0.082t) After 15 years: A = 500 e^(0.082 * 15) = 500 e^1.23 = $1,710.61

OpenStudy (anonymous):

omg thank you!!

OpenStudy (nurali):

My Pleasure.